OpenStudy (anonymous):

Find the solution of 3 square root of x+5=-9

OpenStudy (abmon98):

\[3\sqrt{x+5}=-9\]\[3(x+5)^{1/2}=-9\]\[Expand brackets\]\[3x+15=(-9)^2\]\[3x+15=81\]\[3x=81-15\]\[x=66/3\]

OpenStudy (anonymous):

these are the anser choices: x = 4; not extraneous x = 4; extraneous x = 0; extraneous x = 0; not extraneous